Use the vertical and horizontal algorithms to find the sum or difference of the following.

( Explain the "borrow and carry" concepts.)

a) 587 + 620

b) 749 - 568


1
Expert's answer
2021-05-31T19:28:04-0400

(a) 587+620

Horizontal algorithm-

587=500+80+7

620=600+20+0


587+620= (500+600)+(80+20)+(7+0)=1100+100+7=1207


Vertical Algorithm-


"5^1|8|7\\"

"+6|2|0\\"

"\overline { 12|0|7}"


Here we will add a carry getting after addition of 8 and 2.


(b) 749-568

Horizontal algorithm-

749=700+40+9

568=500+60+8


749-568=(700-500)+(40-60)+(9-8)=200-20+1=181


Vertical algorithm-


"7|4^1|9"

"-5|6|8"

"\overline{ 1|8|1}"


Here we borrow a carry from 7 so that we can perform subtraction.